# finite free probability and $S$ transforms of Jacobi processes

### v1

@@ 5.2.2. Convergence of the finite differences
$\frac{1}{(\ell+1)!}\frac{d^{\ell}}{dz^{\ell+1}}\biggr|_{z=0}G(1/z)=m_{\ell}(\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}).$
Hence, from unicity of analytic continuation $\{G_{\nu^{(d)}_{p_{d}}(v)},\,d\geq 1\}$ has a unique adherence point. Normality implies that $G_{\nu^{(d)}_{p_{d}}(v)}$ converges toward $G_{\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}}$ locally uniformly on $\Omega$. The result is proven.
Case 2: The roots of the sequence $(p_{d})_{g\geq 1}$ are contained in $(0,\eta]$.
Let $v\in(0,1-\mu(\{0\}))$. From Theorem 1.2 in [Ari+24] (and the computations done in Case $1$), $\mu\llbracket\partial^{\lceil{dv}\rceil|d}\rrbracket$ tends to ${\rm Dil}_{v}(\mu^{\boxplus\frac{1}{v}})$. Also, by Lemma 5.4 of [Ari+24], for large enough $j$, $\mu\llbracket\partial^{\lceil{dv}\rceil|d}p_{d}\rrbracket$ has support contained in some $\varepsilon,\eta]$. We apply Case $1$ to infer
$\nabla^{(d)}\mathcal{S}^{(d)}_{p_{d}}(-v)=\frac{d}{\lceil{dv}\rceil}\nabla^{(\lceil{dv}\rceil)}\mathcal{S}^{(\lceil{dv}\rceil)}_{\partial^{\lceil{dv}\rceil|d}p_{d}}(-1)=\frac{1}{v}\partial_{z}S_{{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}}(-1).$ (21)
@@
This case is dealt with as in Section 7.4 in [Ari+24].
@@ 5.3. Finite free $T$-transform of the Frozen Jacobi process
In this section, we assemble the results proved in the last two sections to obtain a differential equation with finite differences satisfied by the finite free $T$ transform of $\chi^{(r,s,m)}_{t}$ of the Hermitian Jacobi process (with parameter $r,s$) and study its high dimensional regime. In the next proposition, $e^{(r,s,m)}_{n}(t)$, $d\geq n\geq 0$ is the $n^{th}$ coefficient of $\chi^{(r,s,m)}(t)$ in the monomial basis $((-1)^{n}x^{d-n}$, $d\geq n\geq 0$ :
$\chi^{(r,s,m)}(t)=\sum_{n=1}^{d}(-1)^{n}e_{n}^{(r,s,m)}(t)x^{d-n}$
@@
$\mathcal{L}_{x}^{(r,s)}\chi_{t}^{(r,s,m)}(x)=$ $x(1-x)\partial_{xx}^{2}+[(r+1)-(r+s+2)x]\partial_{x}\chi_{t}^{r,s,m}$
@@
By equating coefficients on both sides of (22), we obtain
$\frac{d}{dt}e^{(r,s,m)}_{0}(t)$ $=m(2m+r+s-d)e^{(r,s,m)}_{0}(t)=m(p+q-d)e^{(r,s,m)}_{0}(t)=0,$
and for $k\geq 1$,
@@
In the next proposition, we denote by $\lceil{\cdot}\rceil$ the ceiling function.
Proposition 5.3.2.
Let $z\in(0,1)$ and $t\geq 0$. Let $r,s,m$ be as before
$\partial_{t}T_{t}^{(r,s,m)}(z)=[2(m-\lceil{mz}\rceil)-(p+q)]T_{t}^{(r,s,m)}(z)+p-2(m-\lceil{mz}\rceil)\\ +(\frac{\nabla^{(m)}T_{t}^{(r,s,m)}(z)}{T_{t}^{(r,s,m)}(z+\frac{1}{m})})(p-m+\lceil{mz}\rceil+1))({m-\lceil{mz}\rceil})$
@@
$\partial_{t}T^{(r,s,m)}(z)=\frac{n}{m-n+1}\frac{\frac{d}{dt}e^{(r,s,m)}_{n}(t)}{e^{(r,s,m)}_{n-1}(t)}-T_{t}^{(r,s,m)}(z)\frac{\frac{d}{dt}e^{(r,s,m)}_{n-1}(t)}{e^{(r,s,m)}_{n-1}(t)}.$
By using Proposition 5.3.1, we infer the following equation for the derivative of $T_{t}^{(r,s,m)}(z)$:
$\partial_{t}T_{t}^{(r,s,m)}(z)$ $=\frac{n}{m-n+1}\frac{\frac{d}{dt}e^{(r,s,m)}_{n}(t)}{e^{(r,s,m)}_{n-1}(t)}-T_{t}^{(r,s,m)}(z)\frac{\frac{d}{dt}e^{(r,s,m)}_{n-1}(t)}{e^{(r,s,m)}_{n-1}(t)}$
$=(n(n-1)-n(p+q))T_{t}^{(r,s,m)}(z)$
$\hskip 28.45274pt+{n}(p-(n-1))$
$\hskip 56.9055pt-T_{t}^{(r,s,m)}(z)(n-1)((n-2)-(p+q))$
$\hskip 85.35826pt-\frac{T_{t}^{(r,s,m)}(z)}{T_{t}^{(r,s,m)}(z+\frac{1}{m})}(p-(n-2))({n-1})$
We write the quotient:
@@
and replace $n-1=m-\lceil{mv}\rceil$ to obtain the result. ∎
We can now state the main theorem of this section: Theorem 5.3.3 derives the PDE satisfied by the $T$ transform of the free Jacobi process as the limit of the “discrete” PDE stated in Proposition 5.3.2.
@@
We already now that for each $t$, the finite free $T$ transform $T^{(r,s,m)}_{t/d(m)}$ converges. Given Proposition 5.3.2, the only point missing is the locally uniform convergence of the trajectory $\{T^{(r,s,m)}_{t/d(m)},t\geq 0\}$. Let $f:\mathbb{R}\to\mathbb{R}$ be a smooth function with bounded derivative; this will be sufficient to show that
$\mu_{t}^{(r,s,m)}(f),\,t\geq 0$
converges locally uniformly in time. Again, this is readily implied by local equicontinuity. Straightforward computations then yield:
$\partial_{t}\mu^{(r,s,m)}_{t}(f)=\frac{p}{d(m)}+\frac{p+q}{d(m)}\big\{\frac{1}{m}\sum_{j}x_{j}f^{\prime}(x_{j})\big\}+\frac{1}{2md(m)}\sum_{j\neq k}\frac{f(x_{j})-f(x_{k})}{x_{j}-x_{k}}(x_{j}(1-x_{k})+x_{k}(1-x_{j})).$
Further,
@@
Hence, the $\|\partial_{t}\mu_{t}^{(r,s,m)}(f)\|_{\infty}$ is bounded uniformly on $d$ and equicontinuity follows. ∎
Of course, it possible to derive a PDE satisfied by the $S$ transform of the free Jacobi process straightforward computations yield:
$\partial_{t}\mathcal{S}_{t}^{(\lambda,\theta)}(z)=(2\lambda\theta v+1)\mathcal{S}_{t}^{(\lambda,\theta)}(z)-\theta(1+2\lambda z)[\mathcal{S}_{t}^{(\lambda,\theta)}(z)]^{2}+\frac{\theta}{2}z(1+\lambda z)\partial_{z}[\mathcal{S}_{t}^{(\lambda,\theta)}]^{2}(z).$ (23)
This is in accordance with Theorem 3.1.1.
Remark 5.3.4.
Let us emphasize a point of interest with this derivation of the limiting $T$ transform. Suppose we want to study the limiting distribution of a one-parameter family of polynomials with positive roots $(p^{(d)})_{t\geq 0}$, when the degree $d$ goes to infinity. Assume that the coefficients $e_{k}(t),k\leq d$ in the monomial basis of $p^{(d)}_{t}$ satisfy an ODE of the form :
$\frac{d}{dt}e^{(d)}_{k}(t)=\sum_{q=-p}^{p}a^{(k)}_{q}e^{(d)}_{k+q}(t),\quad k\in[p,d-p].$
We assume that the discrete derivatives $\nabla^{(d)}a_{q}^{(\cdot,d)}([zd])=k(a_{q}^{(k,d)}-a_{q}^{(k-1,d)})$, $z\in[\frac{p}{d},1-\frac{p}{d}]$ converge as $d$ goes to infinity and we let $a_{q}(z)$ the limit ($z\in(0,1)$). We make the additional assumptions that the finite free $S$ transforms of $\{p^{(d)}_{t},t\geq 0\}$ converge locally uniformly in the variable $t$ toward $S_{t}(\cdot),\quad t\geq 0$. By definition,
$S^{(d)}_{p^{(d)}_{t}}\left(-\frac{k}{d}\right)=\frac{k}{d-k+1}\frac{e^{(d)}_{k-1}(t)}{e^{(d)}_{k}(t)}.$
Let us set $S^{(d)}_{t}:=S^{(d)}_{p_{t}^{(d)}}$ for readability. Then, straightforward computations yield
$\partial_{t}S^{(d)}(-\frac{k}{d})=-S^{(d)}_{t}\left(-\frac{k}{d}\right)\bigg(a_{0}^{(k,d)}+\sum_{q=1}^{p}a_{q}^{(k,d)}\prod_{1\leq s\leq q}\bigg[S^{(d)}_{t}\left(-\frac{k+s}{d}\right)^{-1}\frac{k+s}{d-(k+s)+1}\bigg]\\ +\sum_{q=1}^{p}a_{-q}^{(k,d)}\prod_{1\leq s\leq q}\bigg[S^{(d)}_{t}\left(-\frac{k-s+1}{d}\right)\frac{d-k+s}{k-s+1}\bigg]\bigg)\\ +\frac{k}{d-k+1}\bigg(a_{1}^{(k-1,d)}+\sum_{q=2}^{p}a_{q}^{(k-1,d)}\prod_{1\leq s\leq q-1}\bigg[S^{(d)}_{t}\left(-\frac{k+s}{d}\right)^{-1}\frac{k+s}{d-(k+s)+1}\bigg]\\ +a_{0}^{(k-1,d)}S^{(d)}_{t}\left(-\frac{k+1}{d}\right)\frac{d+k}{k+1}+\sum_{q=1}^{p}a_{-q}^{(k-1)}\prod_{1\leq s\leq q+1}\bigg[S^{(d)}_{t}\left(-\frac{k-s+1}{d}\right)\frac{d-k+s}{k-s+1}\bigg]\bigg).$
We might go further and write the derivative $\partial_{t}S(-\frac{k}{d})$ as a sum of terms
$a_{0}^{(k-1,d)}\bigg(S^{(d)}_{t}\left(-\frac{k+1}{d}\right)\frac{k(d+k)}{(d-k+1)(k+1)}-S^{(d)}_{t}\left(-\frac{k}{d}\right)\bigg)+S^{(d)}_{t}\left(-\frac{k}{d}\right)(a_{0}^{(k-1,d)}-a_{0}^{(k,d)})$
and
$a_{1}^{(k)}S^{(d)}_{t}\left(-\frac{k}{d}\right)S^{(d)}_{t}\left(-\frac{k+1}{d}\right)^{-1}\frac{k+1}{d-k}+\frac{k}{d-k+1}a_{1}^{(k-1)}=\\ a_{1}^{(k)}\left\{\frac{k}{d-k+1}-S^{(d)}_{t}\left(\frac{k}{d}\right)S^{(d)}_{t}\left(-\frac{k+1}{d}\right)^{-1}\frac{k+1}{d-k}\right\}+\frac{k(a_{1}^{(k-1)}-a_{1}^{(k)})}{d-k+1}$
and for any $q\in[2,p]$ :
$(a_{q}^{(k,d)}-a_{q}^{(k-1,d)})S^{(d)}_{t}\left(-\frac{k}{d}\right)\prod_{1\leq s\leq q}\bigg[S^{(d)}_{t}\left(-\frac{k+s}{d}\right)^{-1}\frac{k+s}{d-(k+s)+1}\bigg]\\ -a_{q}^{(k-1,d)}\bigg(S^{(d)}_{t}\left(-\frac{k}{d}\right)\prod_{1\leq s\leq q}\bigg[S^{(d)}_{t}\left(-\frac{k+s}{d}\right)^{-1}\frac{k+s}{d-(k+s)+1}\bigg]\\ -\frac{k}{d-k+1}\prod_{1\leq s\leq q-1}\bigg[S^{(d)}_{t}\left(-\frac{k+s}{d}\right)^{-1}\frac{k+s}{d-(k+s)+1}\bigg]\bigg).$
Thanks to Theorem 5.2.1, by letting $d$ tends to infinity ($k=[dz])$, the above expression converges to, for any $z\in(0,1)$ and $q\in[2,p]$:
$a_{q}(z)\left(\frac{z}{1-z}-\partial_{z}\left[S_{t}(-z)^{-(q-1)}\frac{z^{q}}{(1-z)^{q}}\right]\right).$
A similar analysis on the terms corresponding to $q\in[-p,-2]$ will provide an expression for the limit as $d$ tends to infinity in terms of the partial derivatives of the limiting $S$ transform $S_{t}$.

### v3

@@ 5.2.2. Convergence of the finite differences
$\frac{1}{(\ell+1)!}\frac{m^{\ell}}{dz^{\ell+1}}\biggr|_{z=0}G(1/z)=m_{\ell}(\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}).$
Hence, from unicity of analytic continuation $\{G_{\nu^{(m)}_{p_{m}}(v)},\,m\geq 1\}$ has a unique adherence point. Normality implies that $G_{\nu^{(m)}_{p_{m}}(v)}$ converges toward $G_{\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}}$ locally uniformly on $\Omega$. Since $0\in\Omega$, we get, locally uniformly on $v$,
$\nabla^{(m)}\mathcal{S}^{(m)}_{p_{m}}(v)=m(\mathcal{S}^{(m)}_{p_{m}}(v+\frac{1}{m})-\mathcal{S}^{(m)}(v))\rightarrow_{m,+\infty}-G_{\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}}(0)=-\partial_{v}G_{{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}}(0)=-\partial_{v}{S}_{\mu}(-v).$
The result is proven.
Case 2: The roots of the sequence $(p_{m})_{g\geq 1}$ are contained in $(0,\eta]$.
Let $v\in(0,1-\mu(\{0\}))$. From Theorem 1.2 in [arizmendi2024s] (and the computations done in Case $1$), $\mu\llbracket\partial^{\lceil{mv}\rceil|m}\rrbracket$ tends to ${\rm Dil}_{v}(\mu^{\boxplus\frac{1}{v}})$. Also, by Lemma 5.4 of [arizmendi2024s], for large enough $j$, $\mu\llbracket\partial^{\lceil{mv}\rceil|m}p_{m}\rrbracket$ has support contained in some $\varepsilon,\eta]$. We apply Case $1$ to infer
$\nabla^{(m)}\mathcal{S}^{(m)}_{p_{m}}(-v)=\frac{m}{\lceil{mv}\rceil}\nabla^{(\lceil{mv}\rceil)}\mathcal{S}^{(\lceil{mv}\rceil)}_{\partial^{\lceil{mv}\rceil|m}p_{m}}(-1)=\frac{1}{v}\partial_{z}S_{{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}}(-1).$ (25)
@@
This case is dealt with as in Section 7.4 in [arizmendi2024s].
@@ 5.3. Finite free $T$-transform of the Frozen Jacobi process
In this paragraph, we put together the results proved in the last two sections to obtain a differential-difference equation satisfied by the finite free $T$ transform of $\chi^{(r,s,m)}_{t}$ of the Hermitian Jacobi process. Once obtained, we pass to the limit therein and derive a PDE satisfied by the $T$ transform of the free Jacobi process.
Let $e^{(r,s,m)}_{n}(t),0\leq n\leq m,$ be the elementary symmetric functions in the roots of $\chi^{(r,s,m)}(t)$:
$\chi^{(r,s,m)}(t)=\sum_{n=0}^{m}(-1)^{n}e_{n}^{(r,s,m)}(t)x^{m-n},\quad e_{0}^{(r,s,m)}(t)=1.$
The following proposition provides a differential system satisfied by $e^{(r,s,m)}_{n}(t),1\leq n\leq m$. In this respect, the reader readily notices that it is simpler than the one derived in [MR4092353], Lemma 3.1., and corresponding to the Jacobi particle system in [1,1][-1,1].
@@
$\mathcal{L}_{x}^{(r,s)}\chi_{t}^{(r,s,m)}(x)=$ $x(1-x)\partial_{xx}^{2}+[(r+1)-(r+s+2)x]\partial_{x}\chi_{t}^{(r,s,m)}$
@@
By equating coefficients on both sides of (26), we obtain for any $k\geq 1$:
@@
Using this proposition, we obtain the following corollary (we recall the notation $\lceil{\cdot}\rceil$ for the ceiling function):
Corollary 5.3.2.
Let $z\in(0,1)$ and $t\geq 0$. Then
$\partial_{t}T_{t}^{(r,s,m)}(z)=[2(m-\lceil{mz}\rceil)-(p+q)]T_{t}^{(r,s,m)}(z)+p-2(m-\lceil{mz}\rceil)\\ +(\frac{\nabla^{(m)}T_{t}^{(r,s,m)}(z)}{T_{t}^{(r,s,m)}(z+\frac{1}{m})})(p-m+\lceil{mz}\rceil+1))({m-\lceil{mz}\rceil}).$
@@
$\partial_{t}T^{(r,s,m)}(z)=\frac{n}{(m-n+1)e^{(r,s,m)}_{n-1}(t)}\frac{de^{(r,s,m)}_{n}}{dt}(t)-\frac{T_{t}^{(r,s,m)}(z)}{e^{(r,s,m)}_{n-1}(t)}\frac{de^{(r,s,m)}_{n-1}}{dt}(t).$
Appealing to Proposition 5.3.1, we infer the following equation for the derivative of $T_{t}^{(r,s,m)}(z)$:
$\partial_{t}T_{t}^{(r,s,m)}(z)$ $=(n(n-1)-n(p+q))T_{t}^{(r,s,m)}(z)+{n}(p-(n-1))$
$\hskip 28.45274pt-T_{t}^{(r,s,m)}(z)(n-1)((n-2)-(p+q))-\frac{T_{t}^{(r,s,m)}(z)}{T_{t}^{(r,s,m)}(z+\frac{1}{m})}(p-(n-2))({n-1}).$
Writing the quotient:
@@
and keeping in mind the relation $n-1=m-\lceil{mv}\rceil$, we get the desired result after straightforward computations. ∎
We can now state the main theorem of this section: Theorem 5.3.3 derives the PDE satisfied by the $T$ transform of the free Jacobi process as the limit of the differential-difference equation stated in Proposition 5.3.2.
@@
We already know that for each $t$, the finite free $T$ transform $T^{(r,s,m)}_{t/d(m)}$ converges. By the virtue of Corollary 5.3.2, the only point missing is the local (in time) uniform convergence (in $m$) of $T^{(r,s,m)}_{t/d(m)},~t\geq 0$. We show that the root distribution $\mu^{(r,s,m)}_{t/d(m)}$ converges weakly, locally uniformly in time : if $f:[0,1]\to\mathbb{R}$ is a smooth function, $\mu_{t/d(m)}^{(r,s,m)}(f),\,t\geq 0$ converges locally uniformly in time. The point-wise convergence in time of the $T$-transform implies that for every $t\geq 0$, $\mu_{t/d(m)}^{(r,s,m)}(f)$ converges (in the regime 2). Let us compute the derivative in time:
$\partial_{t}\mu^{(r,s,m)}_{t}(f)=\frac{p}{d(m)}+\frac{p+q}{d(m)}\big\{\frac{1}{m}\sum_{j}x_{j}f^{\prime}(x_{j})\big\}\\ +\frac{1}{2md(m)}\sum_{j\neq k}\frac{f(x_{j})-f(x_{k})}{x_{j}-x_{k}}(x_{j}(1-x_{k})+x_{k}(1-x_{j})).$
Furthermore, we can bound this derivative uniformly on $m$ by
@@
We infer that $\{t\mapsto\mu_{t/d(m)}^{(r,s,m)}(f),m\geq 0\}$ is equicontinuous and therefore the local uniform convergence of $\mu^{(r,s,m)}_{t/d(m)}(f)$ as $m$ goes to infinity.
Together we the second Dini’s Theorem (since $T_{t/d(m)}^{(r,s,m)}(\cdot)$ is a non-decreasing step-function), we deduce that the right-hand side of the differential-difference equation of Corollary 5.3.2, if normalize by the factor $d(m)$, converges locally uniformly in time, as follows.
Choosing a $f$ polynomial, we infer that all finite free cumulants of $\chi^{(r,s,m)}_{t}$ converge locally uniformly in time. From the formula recalled in equation (21), we also deduce that all the finite free cumulants of the iterated derivatives of $\chi_{t/d(m)}^{(r,s,m)}$ converge locally uniformly in time (see equations (24)) and therefore the weak convergence, locally uniformly in time, of their root distributions.
The formula (20) will imply local uniform convergence in time of the finite free $S$ and $T$ - transform of $\chi_{t}^{(r,s,m)}$, provided that we can bound away from zero the roots of the iterated derivatives of $\chi_{t}^{(r,s,m)}$, locally uniformly in time. This is done using Lemma 5.5 in [arizmendi2024s], since the cut-off $a$ can be chosen locally uniformly in time.
The limit $T^{(\lambda,\theta)}(z)$ is therefore continuously differentiable in time and for each time $t\geq 0$ and $z\in(-1,0)$:
$\partial_{t}T^{(r,s,m)}_{t/d(m)}(z)\rightarrow T^{(\lambda,\theta)}_{t}(z)$
and we can pass to the limit in both sides of the integro-differential equation of Corollary 5.3.2 to obtain the second assertion of the theorem. ∎